In a carnival organized by Youth 4B, competition is conducted. One of the competition is to answer 20 Mathematics question. The prize offered is grand. The total cash prize offered is RM 490. The second runner up will receive half of the cash prize of the first-runner up. The first-runner up will receive half of the cash prize of first place winner. Calculate the cash prize that could be earned by the first place winner. Only first three winners will receive the prize.
Answer by Boreal(15235)

x=winner's prize
(1/2)x= 2nd place
(1/4)x=3rd place
(7/4)x=490
multiply both sides by 4
7x=1960
x=280
First prize is 280 RM ANSWER
First runner up gets 140 RM
Second runner up gets 70 RM